OpenStudy (anonymous):

If sin A sin B - cos A cos B +1 =0 ,prove that 1+cot A tan B = 0

OpenStudy (anonymous):

we just need to prove that sin A cos B - cos A sin B +1 = sin(A+b)

OpenStudy (anonymous):

thats what i can't do!

OpenStudy (anonymous):

take the right side.

OpenStudy (anonymous):

pls elaborate

OpenStudy (anonymous):

recheck the question.

OpenStudy (anonymous):

k it was a cos instyead of sin corrected

OpenStudy (anonymous):

then edit the question .

OpenStudy (anonymous):

did i t already , maybe refres h

OpenStudy (anonymous):

oooh so the question was wrong, I was really confused

OpenStudy (anonymous):

yeah my mistk sry :)

OpenStudy (anonymous):

sin A cos B - cos A cos B +1 =0 ,prove that 1+cot A tan B = 0

OpenStudy (anonymous):

?? see the ques...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

any further progress ..

OpenStudy (anonymous):

the question ?! is there a problem with the question ?!

OpenStudy (anonymous):

nope ques is "If sin A sin B - cos A cos B +1 =0 ,prove that 1+cot A tan B = 0" and nothin else

OpenStudy (anonymous):

maybe the ques in my book is wrong ...

OpenStudy (anonymous):

\[ \sin A \sin B - \cos A \cos B + 1 = 0 \\ \cos (A+B) = 1 \\ \text{which implies} \\ \sin (A+B) = 0 \\ \sin A \cos B + \sin B \cos A = 0 \\ 1 + \cot B \tan A = 0\] but this answer is only valid if \[A + B = 90^\circ\] I think

OpenStudy (anonymous):

how does 2nd line imply 4th one?

OpenStudy (anonymous):

you don't know this formula sin(A+B)=sin(A)cos(B)+sin(B)cos(A) @problemsolver

OpenStudy (anonymous):

i know ..

OpenStudy (anonymous):

but how does sin(A)cos(B)+sin(B)cos(A) = cosA cos B - sinA sinB -1

OpenStudy (anonymous):

k.. i understand you used that in(90-A) =cos(A)..right??

OpenStudy (anonymous):

there's an s before in(90-A)

OpenStudy (anonymous):

A and B are not any real numbers ....they satisfy cos(A+B)=1 ! So the answer of @AugustinusSextus is correct except the last thing ! ! for the 5'th lin===>6th line ! He divided by sin(B)cos(A) both sides ! clear now ?!

OpenStudy (anonymous):

hahaha oops my mistake, but it's still valid tho :p

OpenStudy (anonymous):

can you show me how to type statements with latex (like implies....) !:) please :) @AugustinusSextus

OpenStudy (anonymous):

k.. thanks both of you